solve the system of equations what is the solution set
x+5y=6
x=7-5y

Found 2 solutions by Alan3354, algebrahouse.com:
Answer by Alan3354(69443)

x+5y=6
x=7-5y --> x+5y = 7
--> 6 = 7 No solution
-------------
If you graph them, the lines are parallel.
No intersection, no solution

Answer by algebrahouse.com(1659)

solve the system of equations what is the solution set


x + 5y = 6
x = 7 - 5y

7 - 5y + 5y = 6 {substituted (7 - 5y), in for x, into top equation}
7 = 6 {the variable disappeared and a false statement is the result}
= no solution

Whenever the variable disappears and you get a statement that is never true
then. there is "no solution". In other words, the two lines never meet, or are parallel.

Whenever the variable disappears and you get a statement that is always true
then, there are "infinite solutions". In other words, the two lines are the same,
having infinite intersecting points.
www.algebrahouse.com